# Robustness-Inspired Defense Against Backdoor Attacks on Graph Neural Networks

### v1

@@ Abstract
Graph Neural Networks (GNNs) have achieved promising results in tasks such as node classification and graph classification. However, recent studies reveal that GNNs are vulnerable to backdoor attacks, posing a significant threat to their real-world adoption. Despite initial efforts to defend against specific graph backdoor attacks, there is no work on defending against various types of backdoor attacks where generated triggers have different properties. Hence, we first empirically verify that prediction variance under edge dropping is a crucial indicator for identifying poisoned nodes. With this observation, we propose using random edge dropping to detect backdoors and theoretically show that it can efficiently distinguish poisoned nodes from clean ones. Furthermore, we introduce a novel robust training strategy to efficiently counteract the impact of the triggers. Extensive experiments on real-world datasets show that our framework can effectively identify poisoned nodes, significantly degrade the attack success rate, and maintain clean accuracy when defending against various types of graph backdoor attacks with different properties.
@@ 1 Introduction
Graphs are pervasive in real world such as social networks fan2019graph , molecular graphs mansimov2019molecular , and knowledge graphs liu2022federated . Graph Neural Networks (GNNs) have shown great ability in node representation learning on graphs. Generally, GNNs adopt a message-passing mechanism that updates a node’s representation by iteratively aggregating information from its neighbors. The resulting node representations retain both node attributes and local graph structure information, benefiting various downstream tasks such as node classification hamilton2017inductive ; kipf2016semi ; velickovic2017graph , graph classification xu2018powerful , and link prediction zhang2018link .
@@
Despite their great performance, recent studies Dai_2023 ; xi2021graph ; zhang2021backdoor ; zhang2024rethinking show that GNNs are susceptible to backdoor attacks. Generally, backdoor attacks involve generating and attaching backdoor triggers to a small set of target nodes, which are then assigned a target class. These triggers can be predefined or created by a trigger generator and usually take the form of a node or subgraph. When a GNN model is trained on a backdoored dataset, it learns to associate the presence of the trigger with the target class. Consequently, during inference, the backdoored model will misclassify test nodes with the trigger attached as belonging to the target class, while still maintaining high accuracy on clean nodes without a trigger attached. Backdoor attacks on GNNs are particularly concerning due to their potential impact on critical real-world applications such as healthcare and financial services. For instance, in medical diagnosis networks, an attacker could embed backdoor triggers into the training data, leading the GNN model to misclassify certain medical conditions when the trigger is present.
@@
Thus, graph backdoor attack and defense are attracting increasing attention and several initial efforts have been taken xi2021graph ; zhang2021backdoor ; zhang2024rethinking ; Dai_2023 . For example, the seminal work SBA zhang2021backdoor adopts randomly generated graphs as triggers for graph backdoor attack. To improve the attack success rate, GTA xi2021graph adopts a backdoor trigger generator to generate more powerful sample-specific triggers. Observing that the backdoor trigger and target nodes tend to have low feature similarity, Dai et al. Dai_2023 propose a defense method called Prune, which removes edges that connect nodes with low cosine similarity. It significantly reduces the attack success rate (ASR) of previous works. To improve the stealthiness of backdoor attacks, they proposed unnoticeable backdoor attack method UGBA by maximizing the cosine similarity between backdoor triggers and target nodes. Zhang et. al., zhang2024rethinking show that backdoor triggers tend to be outliers which can be removed with graph outlier detection (OD). To address this, they further proposed DPGBA which can generate in-distribution triggers. Despite initial efforts on backdoor defense, they generally utilize backdoor specific properties to defend against specific backdoor and are ineffective across various types of backdoor triggers and attack methods.
@@
Our main contributions are: (i) We empirically verify that poisoned nodes typically exhibit large prediction variance under edge dropping. (ii) Theoretical analysis guarantees that our specially designed graph convolution operations can precisely distinguish poisoned nodes from clean nodes through random edge dropping. (iii) We propose novel training strategy to train a backdoor robust GNN model even though some poisoned nodes are not identified. (iv) Extensive experiments show the effectiveness of RIGBD in defending against backdoor attacks and maintaining clean accuracy.
@@ 2 Related Work
Graph Backdoor Attacks. SBA zhang2021backdoor is a seminal work on graph backdoor attacks, which adopts randomly generated graphs as triggers. GTA xi2021graph adopts a backdoor trigger generator to generate more powerful sample-specific triggers to improve the attack success rate. UGBA Dai_2023 introduces an unnoticeable loss function aimed at maximizing the cosine similarity between backdoor triggers and target nodes to improve the stealthiness of their attack. DPGBA zhang2024rethinking introduces an outlier detector and uses adversarial learning to generate in-distribution triggers, addressing low ASR or outlier issues in existing graph backdoor attacks. More about graph backdoor attacks are in Appendix A.
Graph Backdoor Defense. UGBA Dai_2023 denotes that the attributes of triggers differ significantly from the attached poisoned nodes in GTA xi2021graph , thereby violating the homophily property typically observed in real-world graphs. Thus they propose a defense method called Prune, which removes edges that connect nodes with low similarity. DPGBA zhang2024rethinking further indicates that although the triggers in UGBA Dai_2023 may demonstrate high similarity to target nodes, the triggers in both UGBA Dai_2023 and GTA xi2021graph are still outliers. Thus they propose a defense method called OD, which involves training a graph auto-encoder and filtering out nodes with high reconstruction loss. More details about backdoor defense in non-structure data li2021anti and robust GNNs zhang2020gnnguard ; zhu2019robust ; wang2021certified are shown in Appendix A. Our work is inherently different from theirs: (i) we aim to defend against various types of graph backdoor attacks; (ii) we design a novel and theoretically guaranteed method to precisely identify poison nodes, aiming to degrade the attack success rate (ASR) without compromising clean accuracy.
@@ 3.2 Prediction Variance Caused by Trigger Edge Dropping
Generally, graph backdoor attacks establish an edge between a trigger and a target node xi2021graph ; Dai_2023 ; zhang2021backdoor , which means the backdoor attack will fail if the edge linking the trigger and the target node is dropped. Furthermore, an explicit requirement for backdoor attacks is that the backdoored GNN model $f_{b}$ behaves normally for clean nodes without triggers attached. Based on the analysis above, we make an assumption: given a backdoored graph $\mathcal{G}_{T}$, if the trigger edge $e_{i}\in\mathcal{E}_{B}$ linking the target node $v_{i}\in\mathcal{V}_{B}$ and trigger $g_{i}$ is dropped, the prediction logits of the backdoored model for the poisoned target node $v_{i}$ will typically change significantly, compared to the prediction change on the nodes connected by clean edges $e_{j}\in\mathcal{E}_{T}\setminus\mathcal{E}_{B}$ when these edges $e_{j}$ are dropped.
@@
To empirically verify our assumption, we conduct the following experiments: Given a backdoored graph, we first pretrain a backdoored model on this graph. For each node, we then iteratively drop each of its neighbors one by one and measure the prediction variance caused by dropping trigger edges $\mathcal{E}_{B}$ and clean edges $\mathcal{E}_{T}\setminus\mathcal{E}_{B}$, respectively. We perform experiments on OGB-arxiv NEURIPS2020_fb60d411 with 565 triggers. The attack method is DPGBA zhang2024rethinking . The model architecture is a 2-layer GCN kipf2016semi . From the results shown in Fig. 1 (a), it is evident that dropping adversarial edge connecting backdoor trigger will result in a much larger prediction variance than dropping clean edges. More results on other datasets can be found in Appendix B.
@@
Thus, an intuitive method for identifying a backdoor trigger or poisoned target node is to examine each edge individually. For instance, in a 2-layer GCN kipf2016semi , we remove one of its neighbors at a time for each node and conduct inference based on the remaining 2-hop neighbors. By observing prediction changes with each neighbor removal, we estimate the likelihood of an edge being linked to the backdoor trigger: greater prediction change indicates a higher probability. However, this method has several issues: (i) Scalability: For graphs with high average degree, this method is computationally expensive. Specifically, for an $L$-layer GCN, the time complexity is $\mathcal{O}(LNd^{L}M(d+M))$, where $d$ is the average degree, $N$ is the number of nodes, and $M$ denotes the feature dimension. The analysis is in Appendix C. As $d$ grows, the time complexity grows exponentially with the power $L$, making it impractical for dense graphs. (ii) Utility: When the backdoor trigger is a subgraph with multiple edges linking it to a target node, examining each edge individually becomes ineffective. As shown in Fig. 1 (b), when two edges linking a trigger and a target node, on the OGB-arxiv dataset, most of the prediction variance caused by dropping trigger edges shows similar values to those caused by dropping clean edges, rendering the intuitive method ineffective.
@@
Vulnerability of Clean Model to Backdoor Triggers. Even if we can remove backdoor triggers from the training dataset, a graph backdoor trigger can still potentially lead to a successful attack as the attacker can craft a backdoor trigger that mimics the neighbor of nodes zhang2024rethinking from the target class. To verify this, we adopt the DPGBA zhang2024rethinking attack method and conduct experiments by removing the backdoor triggers from the training dataset. We then train a 2-layer GCN on this clean dataset and report the attack success rate (ASR) on the test dataset, as shown in Table 1. The Random ASR is calculated by $\frac{1}{C}$ as a reference, where $C$ is the number of classes for each dataset. As shown in the table, even when the model is trained on a clean dataset, it can still react to a backdoor trigger that mimics the neighbor of nodes from the target class. Thus, simply removing the backdoor triggers from the training dataset is insufficient to defend against graph backdoor attacks.
@@ 4.1 Identifying Poisoned Target Nodes by Random Edge Dropping
$\small s(i)={\sum}_{k=1}^{K}\mathrm{KL}(\mathbf{p}_{i},\mathbf{p}_{i}^{k}),$ (2)
@@
$\small\mathbf{H}^{(l+1)}=\sigma\left(\mathbf{\tilde{D}}^{-\frac{1}{2}}\mathbf{A}\mathbf{\tilde{D}}^{-\frac{1}{2}}\mathbf{H}^{(l)}\mathbf{W}^{(l)}\right).$ (3)
where $\mathbf{A}$ is the adjacency matrix without self-connections. In contrast, GCN kipf2016semi obtain node representations by $\mathbf{H}^{(l+1)}=\sigma(\mathbf{\tilde{D}}^{-\frac{1}{2}}(\mathbf{A}+\mathbf{I})\mathbf{\tilde{D}}^{-\frac{1}{2}}\mathbf{H}^{(l)}\mathbf{W}^{(l)})$ where $\mathbf{I}$ is the identity matrix.
We adopt the graph convolution strategy in Eq. (3) to get node representations for the following reasons: (1) For clean nodes, as each neighbor has the same drop ratio $\beta$, in expectation, the node representation is unchanged because the proportion of the expected contribution of each neighbor’s features to the node representation is maintained. Thus, the expectation of prediction variance for clean nodes tends to be small. In contrast, if we include the attributes of the central node, the expectation of the node representation will focus more on the central node after random edge dropping, making the prediction variance unpredictable. We will analyze further in Theorem 1. (2) For poisoned target node, as backdoor triggers are designed to attack various types of targets, once the trigger exists after random edge dropping, it will still lead to a successful attack, regardless of whether the node representations are obtained only based on neighbors or also consider the node itself. However, if the trigger is dropped, the model will exhibit a large prediction variance. Notably, while the attributes of the central node are important for accurate classification, our focus here is on testing the prediction variance and identifying poisoned nodes. Once target nodes are identified, various GNNs can be used to train a classifier with our robust loss in Eq. (6).
@@
Assumptions on Graphs. Following dai2022label ; ma2021homophily , we consider a graph $\mathcal{G}$, where each node $v_{i}$ has features $\mathbf{x}_{i}\in\mathbb{R}^{m}$, a label $y_{i}$ and $\operatorname{deg}(i)$ denotes the number of its neighbors. We assume that: (1) The feature of node $v_{i}$ is sampled from the feature distribution $F_{y_{i}}$ that depends on the label $y_{i}$ of the node, i.e., $\mathbf{x}_{i}\sim\mathcal{F}_{y_{i}}$. (2) Feature dimensions of $\mathbf{x}_{i}$ are independent to each other; (3) The features in $\mathbf{X}$ are bounded by a positive scalar $B$, i.e., $\max_{i,j}|\mathbf{X}[i,j]|\leq B$.
@@
Consider a graph $\mathcal{G}=\left\{\mathcal{V},\mathcal{E},\mathbf{X}\right\}$ following Assumptions (1)-(3). For clean node $v_{i}\in\mathcal{V}$ and its neighbors $\mathcal{N}(i)$, the expectation of the pre-activation output of a single operation defined in Eq. (3) is given by $\mathbb{E}[\mathbf{h}_{i}]$. Then the expectation of the pre-activation output of a single operation defined in Eq. (3) after random edge dropping is given by $\mathbb{E}[\mathbf{h}_{i}^{k}]=\mathbb{E}[\mathbf{h}_{i}]$.
The proof is in Appendix D.1. Theorem 1 shows that, in expectation, the output embedding of a node remains unchanged before and after random edge dropping when applying the layer operation defined in Eq. (3), regardless of the drop ratio. However, in practice, conducting numerous experiments is not feasible. Thus, we analyze the distance between the observed node embeddings after random edge dropping and the expectation of node embeddings without perturbation.
@@
$\small\mathbb{P}\left(\left\|\mathbf{h}_{i}^{k}-\mathbb{E}\left[\mathbf{h}_{i}\right]\right\|_{2}\geq t\right)\leq 2\cdot M\cdot\exp\left(-\frac{\operatorname{deg}(i)_{k}t^{2}}{2\rho^{2}(\mathbf{W})B^{2}M}\right),$ (4)
@@
Based on the above analysis, we theoretically demonstrate that our method, which involves random edge dropping and obtaining node representations through graph convolution as defined in Eq. (3), results in small changes to node representations. Thus, for clean nodes, this stability ensures that the prediction output is consistent as the backdoored model does not exhibit sensitivity towards any particular clean neighbors. However, for poisoned target nodes, the scenario differs. Despite the stability in their node representations after random edge dropping, the last-layer classifier’s sensitivity chen2018detecting to the presence or absence of the backdoor trigger leads to a significant variance in prediction outcomes compared to clean nodes. This is verified in Sec. 3.2.
@@
To empirically verify that our method of random edge dropping can efficiently distinguish poisoned nodes from clean nodes, we adopt the attack method DPGBA and conduct experiments on Cora and OGB-arxiv to show the prediction variance of poisoned nodes and clean nodes after random edge dropping. The details of the experiment settings and the results are in Appendix E. From the results, we observe that: (i) Our method consistently results in higher prediction variance for poisoned nodes, enabling the distinction between poisoned and clean nodes. (ii) Even when two edges link a backdoor trigger to a poisoned node, our method maintains high prediction variance for most poisoned nodes. This showcases the superior performance of our method compared to the intuitive approach of dropping each edge individually. Furthermore, the time complexity of our method scales linearly with $K$, which is typically small, whereas the intuitive method scales exponentially with $d^{L}$, leading to significantly higher computational costs as $d$ increases.

### v2

# Robustness Inspired Graph Backdoor Defense

@@ Abstract
Graph Neural Networks (GNNs) have achieved promising results in tasks such as node classification and graph classification. However, recent studies reveal that GNNs are vulnerable to backdoor attacks, posing a significant threat to their real-world adoption. Despite initial efforts to defend against specific graph backdoor attacks, there is no work on defending against various types of backdoor attacks where generated triggers have different properties. Hence, we first empirically verify that prediction variance under edge dropping is a crucial indicator for identifying poisoned nodes. With this observation, we propose using random edge dropping to detect backdoors and theoretically show that it can efficiently distinguish poisoned nodes from clean ones. Furthermore, we introduce a novel robust training strategy to efficiently counteract the impact of the triggers. Extensive experiments on real-world datasets show that our framework can effectively identify poisoned nodes, significantly degrade the attack success rate, and maintain clean accuracy when defending against various types of graph backdoor attacks with different properties. Our code is available at: github.com/zzwjames/RIGBD.
@@ 1 Introduction
Graphs are pervasive in real world such as social networks (Fan et al., 2019), molecular graphs (Mansimov et al., 2019), and knowledge graphs (Liu et al., 2022). Graph Neural Networks (GNNs) have shown great ability in node representation learning on graphs. Generally, GNNs adopt a message-passing mechanism that updates a node’s representation by iteratively aggregating information from its neighbors. The resulting node representations retain both node attributes and local graph structure information, benefiting various downstream tasks such as node classification (Hamilton et al., 2017; Kipf & Welling, 2016; Velickovic et al., 2017) and graph classification (Xu et al., 2018).
@@
Despite their great performance, recent studies (Dai et al., 2023; Xi et al., 2021; Zhang et al., 2021; 2024) show that GNNs are susceptible to backdoor attacks. Generally, backdoor attacks involve generating and attaching backdoor triggers to a small set of target nodes, which are then assigned a target class. These triggers can be predefined or created by a trigger generator and usually take the form of a node or subgraph. When a GNN model is trained on a backdoored dataset, it learns to associate the presence of the trigger with the target class. Consequently, during inference, the backdoored model will misclassify test nodes with the trigger attached as belonging to the target class, while still maintaining high accuracy on clean nodes without a trigger attached. Backdoor attacks on GNNs are particularly concerning due to their potential impact on critical real-world applications such as healthcare and financial services. For instance, in medical diagnosis networks, an attacker could embed backdoor triggers into the training data, leading the GNN model to misclassify certain medical conditions when the trigger is present.
@@
Thus, graph backdoor attack and defense are attracting increasing attention and several initial efforts have been taken (Xi et al., 2021; Zhang et al., 2021; 2024; Dai et al., 2023). For example, the seminal work SBA (Zhang et al., 2021) adopts randomly generated graphs as triggers for graph backdoor attack. To improve the attack success rate, GTA (Xi et al., 2021) adopts a backdoor trigger generator to generate more powerful sample-specific triggers. Observing that the backdoor trigger and target nodes tend to have low feature similarity, Dai et al. (2023) propose a defense method called Prune, which removes edges that connect nodes with low cosine similarity. It significantly reduces the attack success rate (ASR) of previous works. To improve the stealthiness of backdoor attacks, they proposed unnoticeable backdoor attack method UGBA by maximizing the cosine similarity between backdoor triggers and target nodes. Zhang et al. (2024) show that backdoor triggers tend to be outliers which can be removed with graph outlier detection (OD). To address this, they further proposed DPGBA which can generate in-distribution triggers. Despite initial efforts on backdoor defense, they generally utilize backdoor specific properties to defend against specific backdoor and are ineffective across various types of backdoor triggers and attack methods.
@@
Our main contributions are: (i) We empirically verify that poisoned nodes typically exhibit large prediction variance under edge dropping. (ii) Theoretical analysis guarantees that our specially designed graph convolution operations can precisely distinguish poisoned nodes from clean nodes through random edge dropping. (iii) We propose a novel training strategy to train a backdoor robust GNN model even though some poisoned nodes are not identified. (iv) Extensive experiments show the effectiveness of RIGBD in defending against backdoor attacks and maintaining clean accuracy.
@@ 2 Related Work
Graph Backdoor Attacks. SBA (Zhang et al., 2021) is a seminal work on graph backdoor attacks, which adopts randomly generated graphs as triggers. GTA (Xi et al., 2021) adopts a backdoor trigger generator to generate more powerful sample-specific triggers to improve the attack success rate. UGBA (Dai et al., 2023) introduces an unnoticeable loss function aimed at maximizing the cosine similarity between backdoor triggers and target nodes to improve the stealthiness of their attack. DPGBA (Zhang et al., 2024) introduces an outlier detector and uses adversarial learning to generate in-distribution triggers, addressing low ASR or outlier issues in existing graph backdoor attacks. More about graph backdoor attacks are in Appendix A.
Graph Backdoor Defense. UGBA (Dai et al., 2023) denotes that the attributes of triggers differ significantly from the attached poisoned nodes in GTA (Xi et al., 2021), thereby violating the homophily property typically observed in real-world graphs. Thus they propose a defense method called Prune, which removes edges that connect nodes with low similarity. DPGBA (Zhang et al., 2024) further indicates that although the triggers in UGBA (Dai et al., 2023) may demonstrate high similarity to target nodes, the triggers in both UGBA (Dai et al., 2023) and GTA (Xi et al., 2021) are still outliers. Thus they propose a defense method called OD, which involves training a graph auto-encoder and filtering out nodes with high reconstruction loss. More details about backdoor defense in non-structure data (Li et al., 2021a) and robust GNNs (Zhang & Zitnik, 2020; Zhu et al., 2019; Wang et al., 2021) are shown in Appendix A. Our work is inherently different from theirs: (i) We aim to defend against graph backdoor attacks regardless of the trigger attributes; (ii) we design a novel and theoretically guaranteed method to precisely identify poison nodes, aiming to degrade the attack success rate (ASR) without compromising clean accuracy.
@@ 3.2 Prediction variance caused by trigger edge dropping
Generally, graph backdoor attacks establish an edge between a trigger and a target node (Xi et al., 2021; Dai et al., 2023; Zhang et al., 2021), which means the backdoor attack will fail if the edge linking the trigger and the target node is dropped. Furthermore, an explicit requirement for backdoor attacks is that the backdoored GNN model $f_{b}$ behaves normally for clean nodes without triggers attached. Based on the analysis above, we make an assumption: given a backdoored graph $\mathcal{G}_{T}$, if the trigger edge $e_{i}\in\mathcal{E}_{B}$ linking the target node $v_{i}\in\mathcal{V}_{B}$ and trigger $g_{i}$ is dropped, the prediction logits of the backdoored model for the poisoned target node $v_{i}$ will typically change significantly, compared to the prediction change on the nodes connected by clean edges $e_{j}\in\mathcal{E}_{T}\setminus\mathcal{E}_{B}$ when these edges $e_{j}$ are dropped.
@@
To empirically verify our assumption, we conduct the following experiments: Given a backdoored graph, we first pretrain a backdoored model on this graph. For each node, we then iteratively drop each of its neighbors one by one and measure the prediction variance caused by dropping trigger edges $\mathcal{E}_{B}$ and clean edges $\mathcal{E}_{T}\setminus\mathcal{E}_{B}$, respectively. We perform experiments on OGB-arxiv (Hu et al., 2020) with 565 triggers. The attack method is DPGBA (Zhang et al., 2024). The model architecture is a 2-layer GCN (Kipf & Welling, 2016). From the results shown in Fig. 1 (a), it is evident that dropping adversarial edge connecting backdoor trigger will result in a much larger prediction variance than dropping clean edges. More results on other datasets can be found in Appendix B.
@@
Thus, an intuitive method for identifying a backdoor trigger or poisoned target node is to examine each edge individually. For instance, in a 2-layer GCN (Kipf & Welling, 2016), we remove one of its neighbors at a time for each node and conduct inference based on the remaining 2-hop neighbors. By observing prediction changes with each neighbor removal, we estimate the likelihood of an edge being linked to the backdoor trigger: greater prediction change indicates a higher probability. However, this method has several issues: (i) Scalability: For graphs with high average degree, this method is computationally expensive. Specifically, for an $L$-layer GCN, the time complexity is $\mathcal{O}(LNd^{L}M(d+M))$, where $d$ is the average degree, $N$ is the number of nodes, and $M$ denotes the feature dimension. The analysis is in Appendix C. As $d$ grows, the time complexity grows exponentially with the power $L$, making it impractical for dense graphs. (ii) Utility: When the backdoor trigger is a subgraph with multiple edges linking it to a target node, examining each edge individually becomes ineffective. As shown in Fig. 1 (b), when two edges linking a trigger and a target node, on the OGB-arxiv dataset, most of the prediction variance caused by dropping trigger edges shows similar values to those caused by dropping clean edges, rendering the intuitive method ineffective.
@@
Vulnerability of Clean Model to Backdoor Triggers. Even if we can remove backdoor triggers from the training dataset, a graph backdoor trigger can still potentially lead to a successful attack as the attacker can craft a backdoor trigger that mimics the neighbor of nodes (Zhang et al., 2024) from the target class. To verify this, we adopt the DPGBA (Zhang et al., 2024) attack method and conduct experiments by removing the backdoor triggers from the training dataset. We then train a 2-layer GCN on this clean dataset and report the attack success rate (ASR) on the test dataset, as shown in Table 1. The Random ASR is calculated by $\frac{1}{C}$ as a reference, where $C$ is the number of classes for each dataset. As shown in the table, even when the model is trained on a clean dataset, it can still react to a backdoor trigger that mimics the neighbor of nodes from the target class. Thus, simply removing the backdoor triggers from the training dataset is insufficient to defend against graph backdoor attacks.
@@ 4.1 Identifying poisoned target nodes by random edge dropping
$s(i)={\sum}_{k=1}^{K}\mathrm{KL}(\mathbf{p}_{i},\mathbf{p}_{i}^{k}),$ (2)
@@
$\mathbf{H}^{(l+1)}=\sigma\left(\mathbf{\tilde{D}}^{-\frac{1}{2}}\mathbf{A}\mathbf{\tilde{D}}^{-\frac{1}{2}}\mathbf{H}^{(l)}\mathbf{W}^{(l)}\right).$ (3)
where $\mathbf{A}$ is the adjacency matrix without self-connections. In contrast, GCN (Kipf & Welling, 2016) obtain node representations by $\mathbf{H}^{(l+1)}=\sigma(\mathbf{\tilde{D}}^{-\frac{1}{2}}(\mathbf{A}+\mathbf{I})\mathbf{\tilde{D}}^{-\frac{1}{2}}\mathbf{H}^{(l)}\mathbf{W}^{(l)})$ where $\mathbf{I}$ is the identity matrix.
We adopt the graph convolution strategy in Eq. 3 to get node representations for the following reasons: (1) For clean nodes, as each neighbor has the same drop ratio $\beta$, in expectation, the node representation is unchanged because the proportion of the expected contribution of each neighbor’s features to the node representation is maintained. Thus, the expectation of prediction variance for clean nodes tends to be small. In contrast, if we include the attributes of the central node, the expectation of the node representation will focus more on the central node after random edge dropping, making the prediction variance unpredictable. We will analyze further in Theorem 1. (2) For poisoned target node, as backdoor triggers are designed to attack various types of targets, once the trigger exists after random edge dropping, it will still lead to a successful attack, regardless of whether the node representations are obtained only based on neighbors or also consider the node itself. However, if the trigger is dropped, the model will exhibit a large prediction variance. In cases where all the neighbors of a central node are dropped, we cancel the edge drop operation for that node and retain its original neighbors. Notably, while the attributes of the central node are important for accurate classification, our focus here is on testing the prediction variance and identifying poisoned nodes. Once target nodes are identified, various GNNs can be used to train a classifier with our robust loss in Eq. 6.
@@
Assumptions on Graphs. Following (Dai et al., 2022b; Ma et al., 2021a), we consider a graph $\mathcal{G}$, where each node $v_{i}$ has features $\mathbf{x}_{i}\in\mathbb{R}^{m}$, a label $y_{i}$ and $\operatorname{deg}(i)$ denotes the number of its neighbors. We assume that: (1) The feature of node $v_{i}$ is sampled from the feature distribution $F_{y_{i}}$ that depends on the label $y_{i}$ of the node, i.e., $\mathbf{x}_{i}\sim\mathcal{F}_{y_{i}}$. (2) Feature dimensions of $\mathbf{x}_{i}$ are independent to each other; (3) The features in $\mathbf{X}$ are bounded by a positive scalar $B$, i.e., $\max_{i,j}|\mathbf{X}[i,j]|\leq B$.
@@
Consider a graph $\mathcal{G}=\left\{\mathcal{V},\mathcal{E},\mathbf{X}\right\}$ following Assumptions (1)-(3). For clean node $v_{i}\in\mathcal{V}$ and its neighbors $\mathcal{N}(i)$, the expectation of the pre-activation output of a single operation defined in Eq. 3 is given by $\mathbb{E}[\mathbf{h}_{i}]$. Then the expectation of the pre-activation output of a single operation defined in Eq. 3 after random edge dropping is given by $\mathbb{E}[\mathbf{h}_{i}^{k}]=\mathbb{E}[\mathbf{h}_{i}]$.
The proof is in Appendix D.1. Theorem 1 shows that, in expectation, the output embedding of a node remains unchanged before and after random edge dropping when applying the layer operation defined in Eq. 3, regardless of the drop ratio. However, in practice, conducting numerous experiments is not feasible. Thus, we analyze the distance between the observed node embeddings after random edge dropping and the expectation of node embeddings without perturbation.
@@
$\mathbb{P}\left(\left\|\mathbf{h}_{i}^{k}-\mathbb{E}\left[\mathbf{h}_{i}\right]\right\|_{2}\geq t\right)\leq 2\cdot M\cdot\exp\left(-\frac{\operatorname{deg}(i)_{k}t^{2}}{2\rho^{2}(\mathbf{W})B^{2}M}\right),$ (4)
@@
Based on the above analysis, we theoretically demonstrate that our method, which involves random edge dropping and obtaining node representations through graph convolution as defined in Eq. 3, results in small changes to node representations. Thus, for clean nodes, this stability ensures that the prediction output is consistent as the backdoored model does not exhibit sensitivity towards any particular clean neighbors. However, for poisoned target nodes, the scenario differs. Despite the stability in their node representations after random edge dropping, the last-layer classifier’s sensitivity (Chen et al., 2018) to the presence or absence of the backdoor trigger leads to a significant variance in prediction outcomes compared to clean nodes. This is verified in Sec. 3.2.
@@
To empirically verify that our method of random edge dropping can efficiently distinguish poisoned nodes from clean nodes, we adopt the attack method DPGBA and conduct experiments on Cora and OGB-arxiv to show the prediction variance of poisoned nodes and clean nodes after random edge dropping. The details of the experiment settings and the results are in Appendix E. From the results, we observe that: (i) Our method consistently results in higher prediction variance for poisoned nodes, enabling the distinction between poisoned and clean nodes. (ii) Even when two edges link a backdoor trigger to a poisoned node, our method maintains high prediction variance for most poisoned nodes. This showcases the superior performance of our method compared to the intuitive approach of dropping each edge individually. Furthermore, the time complexity of our method scales linearly with $K$, which is typically small, whereas the intuitive method scales exponentially with $d^{L}$, leading to significantly higher computational costs as $d$ increases. We also compare the prediction variance of low-degree nodes from both homophilic and heterophilic graphs against poisoned nodes in Appendix F.